# Optimizing Fine-Tuning through Advanced Initialization Strategies for Low-Rank Adaptation

### v3

@@ Abstract.
The rapid development of parameter-efficient fine-tuning methods has noticeably improved the efficiency of adapting large language models. Among these, LoRA has gained widespread popularity due to its strong balance of effectiveness and parameter efficiency. However, LoRA relies on initializing two low-rank matrices whose product is zero, which limits its ability to effectively activate and leverage the original model weights—creating a potential bottleneck for optimal performance. To address this limitation, we propose IniLoRA, a novel initialization strategy that initializes the low-rank matrices to closely approximate the original model weights. Experimental results indicate that IniLoRA achieves better performance than LoRA across a range of models and tasks. Additionally, we introduce two variants, IniLoRA-$\alpha$ and IniLoRA-$\beta$, both leveraging distinct initialization methods to enhance performance further. Code is available at https://github.com/hihi885/IniLoRA
@@
$c$s: Computing methodologies Neural networks
@@ 1. Introduction
Parameter-efficient fine-tuning (PEFT) methods, such as LoRA, reduce memory and computational overhead by updating only a subset of model parameters, typically through low-rank adaptations (Hu et al., 2022; Valipour et al., 2023; Pan et al., 2024; Liu et al., 2024). However, research on low-rank matrix initialization remains limited. PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024) explore different strategies for initializing low-rank matrices based on singular value decomposition (SVD). LoftQ (Li et al., 2024) jointly optimizes quantized weights and low-rank matrices to enhance fine-tuning performance in quantized models. Beyond these approaches, we propose several novel strategies for low-rank matrix initialization.
@@ 2.1. Variants of LoRA
Zhang et al. (Zhang et al., 2023) used Singular Value Decomposition (SVD) to prune less important singular values for more efficient updates. Yeh et al. (Yeh et al., 2023) introduced LoHa, which uses Hadamard products to increase weight update ranks while maintaining the original number of trainable parameters, and LoKr, which uses the Kronecker product for the same purpose. Nikdan et al. (Nikdan et al., 2024) proposed RoSA, combining low-rank and sparse components to capture both primary and less prominent update directions. Other approaches include adjusting learning rates for LoRA’s low-rank matrices (Hayou et al., 2024), mimicking full fine-tuning by decomposing weights into magnitude and direction (yang Liu et al., 2024), and proving that gradients are low-rank (Zhao et al., 2024). Ren et al. (Ren et al., 2024) introduced MELoRA, which creates and splices small LoRAs, resembling sparse LoRA methods.

### v4

@@ Abstract.
The rapid development of parameter-efficient fine-tuning methods has noticeably improved the efficiency of adapting large language models. Among these, LoRA has gained widespread popularity due to its strong balance of effectiveness and parameter efficiency. However, LoRA relies on initializing two low-rank matrices whose product is zero, which limits its ability to effectively activate and leverage the original model weights—creating a potential bottleneck for optimal performance. To address this limitation, we propose IniLoRA, a novel initialization strategy that initializes the low-rank matrices to closely approximate the original model weights. Experimental results indicate that IniLoRA achieves better performance than LoRA across a range of models and tasks. Additionally, we introduce two variants, IniLoRA-$\alpha$ and IniLoRA-$\beta$, both leveraging distinct initialization methods to enhance performance further. Code is available at https://github.com/xueyongfu11/IniLoRA
@@
journalyear: 2025copyright: ccconference: 2025 The 9th International Conference on Computer Science and Artificial Intelligence; December 12–15, 2025; Beijing, Chinabooktitle: 2025 The 9th International Conference on Computer Science and Artificial Intelligence (CSAI 2025), December 12–15, 2025, Beijing, Chinadoi: 10.1145/3788149.3788198isbn: 979-8-4007-1962-2/2025/12ccs: Computing methodologies Neural networks
@@ 1. Introduction
Parameter-efficient fine-tuning (PEFT) methods, such as LoRA, reduce memory and computational overhead by updating only a subset of model parameters, typically through low-rank adaptations (Hu et al., 2022; Valipour et al., 2023; Pan et al., 2024; Liu et al., 2024b). However, research on low-rank matrix initialization remains limited. PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024) explore different strategies for initializing low-rank matrices based on singular value decomposition (SVD). LoftQ (Li et al., 2024) jointly optimizes quantized weights and low-rank matrices to enhance fine-tuning performance in quantized models. Beyond these approaches, we propose several novel strategies for low-rank matrix initialization.
@@ 2.1. Variants of LoRA
Zhang et al. (Zhang et al., 2023) used Singular Value Decomposition (SVD) to prune less important singular values for more efficient updates. Yeh et al. (Yeh et al., 2023) introduced LoHa, which uses Hadamard products to increase weight update ranks while maintaining the original number of trainable parameters, and LoKr, which uses the Kronecker product for the same purpose. Nikdan et al. (Nikdan et al., 2024) proposed RoSA, combining low-rank and sparse components to capture both primary and less prominent update directions. Other approaches include adjusting learning rates for LoRA’s low-rank matrices (Hayou et al., 2024), mimicking full fine-tuning by decomposing weights into magnitude and direction (Liu et al., 2024a), and proving that gradients are low-rank (Zhao et al., 2024). Ren et al. (Ren et al., 2024) introduced MELoRA, which creates and splices small LoRAs, resembling sparse LoRA methods.